OpenStudy (anonymous):

The design plan for a swimming pool is a rectangle with length 7 m less than its width. The area is 170m^2. Find the width and length of the swimming pool.

OpenStudy (cathyangs):

17 width. 10 length.

OpenStudy (anonymous):

17, 10

OpenStudy (cathyangs):

the area is x(x+7), which distributes to x^2+7x. you can solve this quadratically, or, like what I did, look at the number in question, 170.

OpenStudy (anonymous):

(lenght)x(width) = 170 (L)(W)=170 L = W-7 so, (W)(W-7) = 170 w^2-7w = 170 solve for W and yu get a positive and negative number. As Length and width cannot be negative the positive number is the right answer